### v1

# On Modular Invariants of Truncated Polynomial Rings in low ranks

Abstract.
We verify the conjectures due to Lewis, Reiner and Stanton about the Hilbert series of the invariant ring of the truncated polynomial ring for all parabolic subgroups up to rank $3$. This is done by constructing an explicit set of generators for each invariant ring in question. We also propose a conjecture concerning the action of the Steenrod algebra and the Dickson algebra on a certain naturally occurring filtration of the invariant ring under the general linear group.
Key words and phrases:
Truncated Polynomial Ring, Modular invariant, Dickson invariants, Steenrod algebra, q-binomial, q-multinomial, finite general linear group, Frobenius power
1991 Mathematics Subject Classification:
Primary 54C40, 14E20; Secondary 46E25, 20C20
1. Introduction
Let $\mathbb{F}_{q}$ denote the finite field of $q$ elements, where $q$ is a power of a fixed prime number $p$. Let $I$ denote the ideal generated by the $m$th Frobenius power $x_{1}^{q^{m}},x_{2}^{q^{m}},\ldots,x_{n}^{q^{m}}$, of the polynomial algebra $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$. We denote by ${\mathcal{Q}}_{m}(n)$ or simply ${\mathcal{Q}}_{m}$, if $n$ is fixed, the quotient ring $S/I$. The general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ acts on $S$ by linear substitutions of variables. Since $I$ is a $G$-invariant ideal, ${\mathcal{Q}}_{m}(n)$ inherits a $G$-action from $S$. This paper discusses the problem of computing invariants of the truncated ring ${\mathcal{Q}}_{m}(n)$ under the action of various parabolic subgroups of $G$.
This problem arises in various contexts, in connection with questions in algebraic combinatorics ([14], [15], [10], [4]), algebraic topology ([1], [7], [9]) and number theory [2]. In particular, taking the limit $m$ to infinity yields results about the cofixed space $S_{G}$ of the polynomial ring under the action of the general linear group. This problem is surprisingly still open in modular invariant theory.
Let $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ be a composition of $n$ and denote by $P(\alpha)$ the corresponding parabolic subgroup. In [10], Lewis, Reiner, and Stanton proposed a marvelous series of conjectures about the Hilbert-Poincaré series of the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ for arbitrary $m$ and any composition $\alpha$ of $n$ in terms of a new combinatorial object called the $(q,t)$-binomial coefficient introduced in [15]:
Conjecture 1.1.
([10, Conjecture 1.5]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the parabolic subgroup $P(\alpha)$ is:
$C_{\alpha,m}(t)=\sum_{\beta,\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\;\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
where $e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})$, $B_{i}=\beta_{1}+\ldots+\beta_{i}$, and for a composition $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ of $d$ with partial sum $\alpha_{1}+\ldots+\alpha_{i}=A_{i}$, the $(q,t)$ binomial coefficient $\begin{bmatrix}d\\ \alpha\end{bmatrix}_{q,t}$ is defined as the quotient
$\begin{bmatrix}d\\ \alpha\end{bmatrix}_{q,t}=\frac{\prod_{j=0}^{d-1}(1-t^{q^{d}-q^{j}})}{\prod_{i=1}^{\ell}\prod_{j=0}^{\alpha_{i}}(1-t^{q^{A_{i}}-q^{A_{i-1}+j}})}.$
In particular, when $\alpha=(n)$, they conjectured that
Conjecture 1.2.
([10, Conjecture 1.2]) The Hilbert-Poincaré series of the ring of invariants of ${\mathcal{Q}}_{m}(n)$ under the action of the general linear group $G=\operatorname{GL}_{n}(\mathbb{F}_{q})$ is:
$C_{n,m}(t)=\sum_{k=0}^{\min(n,m)}t^{(n-k)(q^{m}-q^{k})}\;\begin{bmatrix}m\\ k\end{bmatrix}_{q,t}.$
These conjectures are inspired by properties of the $q$-Catalan and $q$-Fuss Catalan numbers, connecting Hilbert-Poincaré series for certain invariant subspaces with the representation theory of rational Cherednik algebras for Coxeter and complex reflection groups. Some partial results about these two conjectures have been obtained (see Drescher-Shepler [4], Goyal [5] and the original paper [10]). Still, how these combinatorial objects occur in the series remains quite mysterious.
In our previous work [6], we have proved their parabolic conjecture for the minimal parabolic subgroup by constructing an explicit basis. We also refined their conjectures by proposing an explicit $\mathbb{F}_{q}$-basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ as follows:
Conjecture 1.3.
A basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is given inductively by the following formula:
$\mathcal{B}_{m}(\alpha)=\{\delta_{s+1;m}^{\alpha_{1}-s}(f\cdot\Phi^{s}g)\colon 0\leq s\leq\min(\alpha_{1},m),f\in\Delta^{m}_{s},g\in\mathcal{B}_{m-s}(\alpha_{2},\ldots,\alpha_{\ell})\}.$
Here $\Delta^{m}_{s}$ is a certain subspace of the Dickson algebra ([3]) of all $\operatorname{GL}_{s}$-invariant polynomials and $\delta_{a;b}$ is the operator constructed in [6]. The Frobenius-like operator $\Phi$ is given by the rules:
  1. (i)
    $\Phi(\delta_{a;b})=\delta_{a+1;b+1}$;
  2. (ii)
    $\Phi(Q_{r,i})=Q_{r+1,i+1}$ and expand multiplicatively.
The reader of [6] may notice that in this conjecture, we have used the core subspace $\Delta^{m}_{s}$ of the Dickson algebra instead of $\nabla^{m}_{s}$ in the original conjecture. The two candidates can be used interchangeably, but $\Delta^{m}_{s}$ is more convenient for explicit calculation, as we will see later.
In the case $\alpha=(n)$ so that $P(\alpha)=G$ is the full general linear group, we have
Conjecture 1.4.
The set $\mathcal{B}_{m}(n)$ consisting of the following family of elements:
$\delta_{s+1;m}^{n-s}(f),f\in\Delta^{m}_{s},0\leq s\leq\min(m,n),$
forms a basis for the $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{G}(n)$.
It is a simple counting problem that our proposed bases have the correct series as predicted. Hence our conjectures refine the two conjectures proposed by Lewis, Reiner, and Stanton. Furthermore, they also explain where the summands of the Hilbert-Poincaré series occur in the original conjectures: The summands of $C_{n,m}(t)$ are just the Hilbert-Poincaré series for $\Delta^{m}_{s}$, $0\leq s\leq\min(m,n)$, shifted by an appropriate degree because of the iterated delta operator.
In [6], we have verified the refined parabolic conjecture 1.3 for the Borel subgroup, that is when $\alpha=1^{n}$. The aim of this note is to verify our conjectures for other parabolic subgroups, including the full general linear group, for $n\leq 3$. As an immediate corollary, the parabolic conjecture of Lewis, Reiner and Stanton is true in these cases.
Theorem 1.5.
The refined parabolic conjectures 1.3 and 1.4 are true for $n\leq 3$ and all $m\geq 1$. Hence, the parabolic conjecture of Lewis, Reiner, and Stanton is true in these cases.
In this paper, we choose to work as explicitly as possible to demonstrate the underlying mechanism of our construction and the additional structure of our basis. The computation reveals an interesting property of the invariant rings considered as a module over the Dickson algebra. We also describe the action of the Steenrod algebra. In subsequent work, we hope to return to this problem in greater generality.
When $n=2$, the only parabolic subgroup other than the Borel subgroup is $\operatorname{GL}_{2}$ itself, and the parabolic conjecture has already been verified in the original paper [10]. However, the authors worked with the dual situation for cofixed spaces and used a particular property of the bivariate case to avoid the truncated ring. Thus, it is unclear whether the invariant subspace ${\mathcal{Q}}_{m}^{G}$, even in rank $2$, can be described explicitly using their method.
In [5, Proposition 4.1], Goyal constructed a family of $G$-invariant polynomials in ${\mathcal{Q}}_{m}(2)$, which he denoted as $y_{k^{\prime}}$ ($0\leq k^{\prime}\leq\frac{q^{m}-q}{q-1}$). It turns out that this family coincides with the $G$-invariant family $\delta_{2}(\Delta^{m}_{1})$ in our notation.
To simplify the notation, we will often drop the index $n$ in ${\mathcal{Q}}_{m}(n)$ when the number of variables is well-understood from the context. Moreover, since the Frobenius power $m$ is fixed, $\delta_{s}$ will be always used in place of $\delta_{s;m}$ throughout the paper. We will also write $(i_{1},i_{2})\in\Delta^{m}_{2}$ instead of $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\in\Delta^{m}_{2}$. Finally we denote by $[a]_{q}$ the $q$-integer $(q^{a}-1)/(q-1)$.
Here is our first result in rank $2$.
Proposition 1.6.
The set $\mathcal{B}_{m}(2)$ consisting of 3 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{G}(2)$ of the full general linear group.
  1. (1)
    $\delta_{1}(\delta_{1}(1))=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}(\Delta^{m}_{1})$.
  3. (3)
    $\Delta^{m}_{2}$.
In rank 3, other than the Borel subgroups, there are 2 proper parabolic subgroups. Their corresponding rings of invariants are described in the next 2 propositions.
Proposition 1.7.
The set $\mathcal{B}_{m}(2,1)$ consisting of 6 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{P(2,1)}$ of the parabolic subgroup corresponding to the composition $(2,1)$.
  1. (1)
    $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}Q_{3,2}^{i}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$, $i<[m-2]_{q}$.
  2. (2)
    $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\delta_{3}(1)$, $(i_{1},i_{2})\in\Delta^{m}_{2}$.
  3. (3)
    $\delta_{2}(Q_{1,0}^{i_{1}}Q_{2,1}^{i})$, $i_{1}<[m]_{q}$, $i<[m-1]_{q}$.
  4. (4)
    $\delta_{2}(Q_{1,0}^{i_{1}}\delta_{2}(1))$, $i_{1}<[m]_{q}$.
  5. (5)
    $\delta_{1}(\delta_{1}(Q_{1,0}^{i}))$, $i<[m]_{q}$.
  6. (6)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))$.
Proposition 1.8.
The set $\mathcal{B}_{m}(1,2)$ consisting of 6 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}^{P(1,2)}$ of the parabolic subgroup corresponding to the composition $(1,2)$.
  1. (1)
    $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$, $j_{1}<[m]_{q},(i_{1},i_{2})\in\Delta^{m-1}_{2}$.
  2. (2)
    $Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{1}<[m]_{q},j_{2}<[m-1]_{q}$.
  3. (3)
    $Q_{1,0}^{j_{1}}\delta_{2}(\delta_{2}(1))$, $j_{1}<[m]_{q}$.
  4. (4)
    $\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$, $(i_{1},i_{2})\in\Delta^{m}_{2}$.
  5. (5)
    $\delta_{1}(\delta_{2}(Q_{1,0}^{j_{1}}))$, $j_{1}<[m]_{q}$.
  6. (6)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))$.
Our final result, which is also the most technical, is for the full general linear group in rank 3:
Proposition 1.9.
The set $\mathcal{B}_{m}(3)$ consisting of 4 families below forms a basis for the invariant subspace ${\mathcal{Q}}_{m}(3)^{G}$ of the full general linear group.
  1. (1)
    $\delta_{1}(\delta_{1}(\delta_{1}(1)))=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}(\delta_{2}(\Delta^{m}_{1}))$.
  3. (3)
    $\delta_{3}(\Delta^{m}_{2})$.
  4. (4)
    $\Delta^{m}_{3}$.
Since the case of the Borel subgroup, which corresponds to the composition $(1,1,1)$ of $n=3$, has been discussed in full generality in [6], the Propositions above combined imply the main Theorem 1.5.
Note that when $m$ is small compared to $n$, all results remain valid with proper interpretation. For example if $n=3$ and $m=2$, then $\Delta^{2}_{3}$ is empty, and $\delta_{3;2}(\Delta^{2}_{2})$ just consists of a single element, which is $1$ in degree 00. To avoid too many special cases (whose proof are not complicated), we will generally assume that $m$ is sufficiently large, comparing to $n$.
In the course of proving these results, we have also computed the structure of these invariant rings, considered as a module over the Dickson algebra, ${\mathcal{D}}_{n}$, and a module over the Steenrod algebra $\mathcal{A}$. At least in the case we have computed, our basis yields a remarkable filtration, namely the subspaces $\mathcal{F}_{n,k}$, for $0\leq k\leq n$, of ${\mathcal{Q}}_{m}^{G}(n)$ spanned by $\delta_{s+1}^{n-s}(\Delta^{m}_{s})$ for all $s\leq k$. We have the following conjecture about this filtration:
Conjecture 1.10.
For each $1\leq k\leq\min(m,n)$, the subspace $\mathcal{F}_{n,k}$ given by
$\mathcal{F}_{n,k}=\operatorname{Span}\{\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq k\},$
is an $\mathcal{A}$-submodule as well as a ${\mathcal{D}}_{n}$-submodule of ${\mathcal{Q}}_{m}^{G}(n)$. Moreover, $\mathcal{F}_{n,k}$ is annihilated by the Dickson invariants $Q_{n,0},Q_{n,1}\ldots Q_{n,n-k-1}$.
This conjecture suggests that the filtration $\mathcal{F}_{n,*}$ may have another description of topological nature. We can prove the following:
Theorem 1.11.
The conjecture 1.10 is true for $n\leq 3$.
The layout of this paper is as follows: Section 2 recalls basic results about modular invariant theory and the key construction from our previous work.
In Section 3, we give a lower bound on the total dimension of the invariant subspaces. The following four Sections are devoted to demonstrating the four Propositions from 1.6 to 1.9 above. The last section discusses the action of the Dickson algebra and the Steenrod algebra and the proof of Theorem 1.11.
Acknowledgement
This work was partially completed while the second and third authors visited the Vietnam Institute for Advanced Study in Mathematics (VIASM). We want to thank Vo Thanh Tung for some computer calculations in the initial stage of this project.
2. Recollection
In this section, we recall some background materials about modular invariant theory and the main construction in [6].
2.1. Upper triangular invariants and Dickson invariants
The canonical projection $S\to{\mathcal{Q}}_{m}$ is clearly a $\operatorname{GL}$-map, hence for each composition $\alpha$ of $n$, there is an induced $\mathbb{F}_{q}$-algebra homomorphism
$S^{G}\subset S^{P(\alpha)}\to{\mathcal{Q}}_{m}^{P(\alpha)},$
making the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}$ into a module over $S^{G}$. Thanks to the work of Dickson [3] since 1911, the structure of the algebra $S^{G}$, which we now call the Dickson algebra, is well-known, and plays a significant role in our investigation.
For each positive integer $k$, denote by $V_{k}$ the product
$V_{k}(x_{1},\ldots,x_{k})=\prod_{\lambda_{i}\in\mathbb{F}_{q}}(x_{k}+\lambda_{1}x_{1}+\ldots+\lambda_{k-1}x_{k-1}).$
It is well-known that the invariant ring $S^{B}$ under the Borel subgroup of upper triangular matrices is a polynomial algebra generated by $V_{i}^{q-1}$, $1\leq i\leq n$. Let $X$ be an indeterminate and recall the fundamental equation
$V_{n+1}(x_{1},\ldots,x_{n},X)=X^{q^{n}}+\sum_{i=0}^{n-1}(-1)^{n-i}Q_{n,i}X^{q^{i}}.$
The polynomials $Q_{n,i}=Q_{n,i}(x_{1},\ldots,x_{n})$ are evidently $\operatorname{GL}_{n}$-invariant. They are called the Dickson invariants, because of the following celebrated theorem due to Dickson:
${\mathcal{D}}_{n}=S^{G}=\mathbb{F}_{q}[Q_{n,0},\ldots,Q_{n,n-1}].$
The Dickson invariants $Q_{n,i}$ can be described in a more explicit form as follows. For any $n$-tuple of non-negative integers $(r_{1},\dots,r_{n})$, put
$[r_{1},\dots,r_{n}]:=\det(x_{i}^{q^{r_{j}}})_{1\leq i,j\leq n}$
and let $L_{n}=L_{n}(x_{1},\ldots,x_{n})$ denote $[0,1\ldots,n-1]$. We then have for $1\leq i\leq n$
$Q_{n,i}=[0,\ldots,\hat{i},\ldots,n]/L_{n},$
In particular, $Q_{n,n}=1$ and $Q_{n,0}=L_{n}^{q-1}$. By convention, we set $Q_{n,i}=0$ if $i<0$ or $i>n$. The following relations are well-known (see for example [13]):
$L_{n}=V_{1}\cdots V_{n},\quad Q_{n,i}=Q_{n-1,i-1}^{q}+Q_{n-1,i}V_{n}^{q-1}.$
Note that $L_{n}$ is, up to sign, the product of all lines in the $\mathbb{F}_{q}$-vector space spanned by $x_{1},\ldots,x_{n}$.
2.2. The $\Delta$ subspace of the Dickson algebra
Given a positive integer $s$ and a partition $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{s}$ with at most $s$ nonzero parts, we say that a Dickson monomial $Q_{s,s-1}^{e_{1}}\ldots Q_{s,0}^{e_{s}}$ in ${\mathcal{D}}_{s}$ is of type $(\lambda_{1},\ldots,\lambda_{s})$ (or q-compatible with $(\lambda_{1},\ldots,\lambda_{s})$, in the language of [15, Definition 5.4]) if for each $1\leq i\leq s$,
$e_{i}\in\bigg{[}\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1}\bigg{)}.$
It is clear that each Dickson monomial is $q$-compatible with a unique partition $(\lambda_{1},\ldots,\lambda_{s})$ of length $\leq s$.
Now if $s\leq m$ then we restrict our attention to partitions which fit inside a rectangular $s\times(m-s)$, that is partitions with at most $s$ nonzero parts and $m-s\geq\lambda_{1}$. Define $\Delta_{(\lambda_{1},\ldots,\lambda_{s})}$ as the subset of the Dickson algebra ${\mathcal{D}}_{s}$ consisting of Dickson monomials of type $(\lambda_{1},\ldots,\lambda_{s})$, and let $\Delta^{m}_{s}$ denote the disjoint union of all subsets over all such partitions fitting inside $s\times(m-s)$. By convention, we put $\Delta^{m}_{s}=\emptyset$ if $s>m$, $\Delta^{m}_{0}=\{1\}$.
The Dickson monomials in $\Delta^{m}_{s}$ for $s\leq\min(m,n)$ together with the delta operators recalled below are the building blocks for our basis. For this reason, we will say that a Dickson monomial in $\Delta^{m}_{s}$ an essential monomial. We will also work with the edge monomials, which is a Dickson monomial not in $\Delta^{m}_{s}$ but has the form $Q_{s,i}f$ where $f\in\Delta^{m}_{s}$, for some $10\leq i\leq s-$ 1. Clearly, the edge monomials are those such that $e_{i}\in[\frac{q^{\lambda_{i}}-q^{\lambda_{i+1}}}{q-1},\frac{q^{\lambda_{i}+1}-q^{\lambda_{i+1}}}{q-1})$ for all $i\neq j$ and $e_{j}=\frac{q^{\lambda_{j}+1}-q^{\lambda_{j+1}}}{q-1}$ for some $1\leq j\leq s$.
It is easy to check that $\Phi\Delta^{m}_{s}\subset\Delta^{m}_{s+1}$. In fact, $\Phi$ sends a Dickson monomial in ${\mathcal{D}}_{s}$ of type $(\lambda_{1},\ldots,\lambda_{s})$ to a Dickson monomial in ${\mathcal{D}}_{s+1}$ of type $(\lambda_{1},\ldots,\lambda_{s},0)$.
2.3. The delta operator
The operator $\delta_{a;b}$ introduced in [6] is fundamental to our construction. We now recall its definition and main properties.
Definition 2.1.
Let $a,b,c$ be positive integers such that $11\leq a\leq c+$ 1. If $f$ is a rational function in $c$ variables, then $\delta_{a;b}(f)$ is a function in $(c+1)$ variables, defined as the quotient
$\delta_{a;b}(f)=\frac{\begin{vmatrix}x_{1}&\cdots&x_{a}\\ x_{1}^{q}&\cdots&x_{a}^{q}\\ \vdots&\ddots&\vdots\\ x_{1}^{q^{a-2}}&\cdots&x_{a}^{q^{a-2}}\\ x_{1}^{q^{b}}f(\widehat{x_{1}},x_{2},\ldots,x_{c+1})&\cdots&x_{a}^{q^{b}}f(x_{1},\ldots,\widehat{x_{a}},\ldots,x_{c+1})\end{vmatrix}}{\begin{vmatrix}x_{1}&\cdots&x_{a}\\ x_{1}^{q}&\cdots&x_{a}^{q}\\ \vdots&\ddots&\vdots\\ x_{1}^{q^{a-1}}&\cdots&x_{a}^{q^{a-1}}\end{vmatrix}}.$
Thus if $b\geq a-1$, $\delta_{a;b}$ increases degree by $q^{b}-q^{a-1}$. By definition, if we separate the first $(a-1)$ variables and write $f$ in the form $f=\sum g(x_{1},\ldots,x_{a-1})h(x_{a},\ldots,x_{c})$, then
$\delta_{a;b}(f)=\sum\delta_{a;b}(g)h(x_{a+1},\ldots,x_{c+1}).$
In this paper, with the number $m$ fixed from the beginning, we will always write $\delta_{s}$ for $\delta_{s;m}$ and occasionally use the full notation $\delta_{a;b}$ when necessary.
It was shown in greater generality in [6] that $\delta_{s}(f)$ is generally not a polynomial, but it is very often a polynomial and is a $\operatorname{GL}_{s}$-invariant. We give a precise statement and provide an elementary proof in the case for $\delta_{3}$ to make the paper self-contained.
Proposition 2.2.
If $f$ is a $\operatorname{GL}_{2}$-invariant polynomial, then $\delta_{3}(f)$ is a polynomial and is $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}(3)$.
Proof.
We first show that $\delta_{3}(f)$ is a polynomial. The denominator $L_{3}$ is a product of linear forms, so it suffices to prove that the numerator becomes zero whenever there exists a nontrivial linear relation $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$. This is reduced to the proof that
$a_{1}x_{1}^{q^{m}}f(x_{2},x_{3})+a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=0,$
provided $a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}=0$ and not all scalars $a_{i}$ are zero. Without loss of generality, assume that $a_{3}=-1$. If both $a_{1},a_{2}\neq 0$, then since $f$ is a $\operatorname{GL}_{2}$-invariant, we have
$f(x_{2},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},a_{1}x_{1}+a_{2}x_{2})=f(x_{1},x_{2}),$
and the above expression becomes
$[a_{1}x_{1}^{q^{m}}+a_{2}x_{2}^{q^{m}}+a_{3}x_{3}^{q^{m}}]f(x_{1},x_{2})=[a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}]^{q^{m}}f(x_{1},x_{2})=0.$
If one of the $a_{i}$ is zero, say $a_{1}=0$, then $f(x_{2},x_{3})=f(x_{2},a_{3}x_{2})=0$ and $f(x_{1},x_{3})=f(x_{1},x_{2})$, so that the sum is reduced to
$a_{2}x_{2}^{q^{m}}f(x_{1},x_{3})+a_{3}x_{3}^{q^{m}}f(x_{1},x_{2})=(a_{2}x_{2}+a_{3}x_{3})^{q^{m}}f(x_{1},x_{2})=0.$
If $a_{1}=a_{2}=0$ or $x_{3}=0$, this is also clear. We have proved that $\delta_{3}(f)$ is a polynomial.
Next we show that $\delta_{3}(f)$ is $\operatorname{GL}_{3}$-invariant modulo $I_{m}$. Clearly, it is symmetric, so it suffices to prove that $\delta_{3}(f)$ does not change under the operation that sends $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$. The denominator $L_{3}$ certainly satisfies this condition. By direct inspection, $\delta_{3}(f)(x_{1}+x_{2},x_{2},x_{3})-\delta_{3}(f)(x_{1},x_{2},x_{3})$ equals
$x_{2}^{q^{m}}\times\frac{\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ f(x_{2},x_{3})-f(x_{1}+x_{2},x_{3})&f(x_{1}+x_{2},x_{3})-f(x_{1},x_{3})&0\end{vmatrix}}{L_{3}}.$
For this expression to be zero in ${\mathcal{Q}}_{m}(3)$, we must prove that the second factor is a polynomial. The rest of the argument is similar to the first part of the proof. ∎
Example 2.3.
The $\operatorname{GL}_{1}$-polynomials are of the form $f(x)=x^{s(q-1)}=Q_{1,0}^{s}$, $s\geq 0$. Then $\delta_{2}(x^{s(q-1)})$ is a $\operatorname{GL}_{2}$-invariant polynomial in ${\mathcal{Q}}_{m}(2)$. For $s<[m]_{q}$, this is Goyal’s $y_{s}$ family in [5] where it was defined as
$y_{s}=x_{1}^{q^{m}-q}x_{2}^{s(q-1)}+x_{1}^{q^{m}-q-(q-1)}x_{2}^{(s+1)(q-1)}+\ldots+x_{1}^{s(q-1)}x_{2}^{q^{m}-q}.$
Our construction is also valid for $s\geq[m]_{q}$. However, it is trivial in ${\mathcal{Q}}_{m}$ except when $s=[m]_{q}+1$, in which case
$\delta_{2}(Q_{1,0}^{[m]_{q}+1})=-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}=-\delta_{1}^{2}(1).$
It is easy to check that $y_{s}$ is a genuine $\operatorname{GL}_{2}$-invariant (that is, a Dickson polynomial) whenever $s=\frac{q^{i}-1}{q-1}$ for some non-negative integer $i$.
Example 2.4.
For $s\leq[m]_{q}$, $\delta_{2}^{2}(Q_{1,0}^{s})$ is a $\operatorname{GL}_{3}$-invariant of degree $2(q^{m}-q)+s(q-1)$ in ${\mathcal{Q}}_{m}$. This family of elements coincides with Goyal’s polynomials $a_{m,3,s}$ in [5, Corollary 4.3] where it was written in a different form:
$a_{m,3,s}=\sum_{(i_{1},i_{2},i_{3})}x_{1}^{i_{1}(q-1)}x_{2}^{i_{2}(q-1)}x_{3}^{i_{3}(q-1)},$
where the sum runs over all triple $(i_{1},i_{2},i_{3})$ of non-negative integers such that $i_{s}<[m]_{q}$ and $i_{1}+i_{2}+i_{3}=2[m]_{q}-2+s$. Again, our definition works also for $s>[m]_{q}$:
$\delta_{2}^{2}(Q_{1,0}^{[m]_{q}+2})=\delta_{1}^{3}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
and is trivial in all other cases.
2.4. The delta operator and the Dickson algebra
For each composition $\alpha$ of $n$, the invariant ring ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is a module over the Dickson algebra ${\mathcal{D}}_{n}$. The next result describes how the delta operator interacts with the Dickson algebra in low rank.
Proposition 2.5.
We have the following identities in ${\mathcal{Q}}_{m}$.
  1. (1)
    $Q_{s,0}\delta_{s}(f)=0$ for all $f$.
  2. (2)
    $Q_{2,1}\delta_{2}(f)=\delta_{2}(Q_{1,0}^{q}f)$ for all $f\in{\mathcal{D}}_{1}$.
  3. (3)
    $Q_{3,i}\delta_{3}(f)=\delta_{3}(Q_{2,i-1}^{q}f)$ for $i=1$ or $2$ and for all $f\in{\mathcal{D}}_{2}$.
  4. (4)
    $Q_{3,2}\delta_{2}^{2}(f)=\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ for all $f\in{\mathcal{D}}_{1}$.
  5. (5)
    $Q_{3,1}\delta_{2}^{2}(f)=0$ for all $f\in{\mathcal{D}}_{1}$.
Proof.
We will prove only statements (3) for $i=1$ and the last two parts, as the proof of other statements use similar arguments and are even more straightforward.
Proof of (3). Consider the difference $Q_{3,1}\delta_{3}(f)-\delta_{3}(Q_{2,0}^{q}f)$. Using the identity $Q_{3,1}=V_{3}^{q-1}Q_{2,1}+Q_{2,0}^{q}$, the (3,1)-entry in the last row of the determinant in the numerator can be simplified as
$x_{1}^{q^{m}}V_{3}^{q-1}(x_{2},x_{3},x_{1})Q_{2,1}(x_{2},x_{3}).$
Since $V_{3}(x_{2},x_{3},x_{1})L_{2}(x_{2},x_{3})=L_{3}(x_{2},x_{3},x_{1})=L_{3}(x_{1},x_{2},x_{3})$, the Laplace expansion along the last row of the determinant in the numerator shows that the resulting polynomial belongs to $I_{m}$, which is zero in ${\mathcal{Q}}_{m}$.
Proof of (4). Write $Q_{3,2}=V_{3}^{q-1}+Q_{2,1}^{q}$. From (1) and (2) we have in ${\mathcal{Q}}_{m}$
$Q_{3,2}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,1}^{q}\delta_{2}f).$
Using the formula $Q_{2,1}=V_{2}^{(q-1)}+Q_{1,0}^{q}$ and unravelling the definition, we see that the difference $\delta_{2}(Q_{2,1}^{q}\delta_{2}f)-\delta_{2}^{2}(Q_{1,0}^{q^{2}}f)$ can be written as the sum
$x_{3}^{q^{m}}\frac{x_{1}^{q^{m}}x_{2}f(x_{2})V_{2}^{q^{2}-q-1}(x_{2},x_{3})-x_{2}^{q^{m}}x_{1}f(x_{1})V_{2}^{q^{2}-q-1}(x_{1},x_{3})}{L_{2}(x_{1},x_{2})}+\\ x_{1}^{q^{m}}x_{2}^{q^{m}}f(x_{3})\frac{x_{2}V_{2}^{q^{2}-q-1}(x_{3},x_{2})-x_{1}V_{2}^{q^{2}-q-1}(x_{3},x_{1})}{L_{2}(x_{1},x_{2})}.$
It is clear that the first summand is a polynomial, and is a multiple of $x_{3}^{q^{m}}$. The second summand belongs to the ideal $(x_{1}^{q^{m}},x_{2}^{q^{m}})$.
Proof of (5). We have in ${\mathcal{Q}}_{m}$
$Q_{3,1}\delta_{2}(\delta_{2}f))=\delta_{2}(Q_{2,0}^{q}\delta_{2}f)$
and
$Q_{2,0}^{q}\delta_{2}(f)=L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{1}f(x_{1})x_{2}^{q^{m}}-L_{2}^{q^{2}-q-1}(x_{1},x_{2})x_{2}f(x_{2})x_{1}^{q^{m}}.$
The rest of the argument is similar to (4) above. ∎
2.5. The Borel subgroup
In [6], we have constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}^{B}$ under the action of the Borel subgroup for arbirary $m$ and $n$. In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists 4 families:
  1. (1)
    $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{j_{3}(q-1)}$, $j_{3}\leq[m]_{q}$.
  2. (2)
    $x_{1}^{q^{m}-1}x_{2}^{j_{2}(q-1)}Q_{2,1}(x_{2},x_{3})^{j_{3}}$, $j_{2}<[m]_{q},j_{3}\leq[m-1]_{q}$.
  3. (3)
    $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$, $j_{1}<[m]_{q}$, $j_{2}\leq[m-1]_{q}$.
  4. (4)
    $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$, $j_{1}<[m]_{q}$, $j_{2}<[m-1]_{q}$, $j_{3}\leq[m-2]_{q}$.
3. A lower bound on the total dimension of the invariant subspaces
The original parabolic conjecture predicted that the Hilbert-Poincaré series for the space of invariants ${\mathcal{Q}}_{m}^{\alpha}(n)$, where $\alpha$ is a composition of $n$, is the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is given in 1.1. In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ should be the total dimension of the graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$.
In this section, we will show that:
Proposition 3.1.
For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$ is at least $C_{\alpha,m}(1)$.
This result is already implicit in the discussion in Section 6 of [10], and is valid for any $n$. This relatively simple remark has the following important consequence in practice: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ such that, by dimension counting, the Hilbert-Poincare series $C^{\prime}_{m,\alpha}(t)$ for ${\mathcal{Q}}_{m}^{\alpha}(n)$ is not greater than $C_{\alpha,m}(t)$ ($f(t)\leq g(t)$ iff $g-f$ is a polynomial with non-negative coefficients). Then since there is a reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ when evaluated at $t=1$, we can conclude that these two series are identical since both are finite degree polynomials with non-negative coefficients. Hence our generating set is actually a basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$.
Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra
$\mathcal{R}=\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]/(x_{1}^{q^{m}}-x_{1},\ldots,x_{n}^{q^{m}}-x_{n}).$
The action of $\operatorname{GL}_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
$\{0\}=F_{-1}\subset F_{0}\subset F_{1}\subset\cdots\subset F_{n(q^{m}-1)}={\mathcal{R}},$
where $F_{i}$ is the image of the set of all polynomials of degree at most $i$ under the canonical projection $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]\twoheadrightarrow{\mathcal{R}}$. Let $\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}$ denote the associated graded vector space $\bigoplus_{i\geq 0}F_{i}/F_{i-1}$ and equip $\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}$ with a ring structure where the multiplication is induced by the product $F_{i}F_{j}\to F_{i+j}$. The canonical $\operatorname{\mathbb{F}}_{q}$-algebra map
$\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]\to{\mathcal{R}}\to\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}},\quad x_{i}\mapsto\bar{x}_{i}\in F_{1}/F_{0},$
then induces an isomorphism of $\operatorname{GL}_{n}$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
${\mathcal{Q}}_{m}(n)\cong\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}}.$
On the other hand, if we apply the functor taking $\operatorname{GL}_{n}$-invariant on each short exact sequence of $\operatorname{GL}_{n}$-modules
$0\to F_{i-1}\to F_{i}\to F_{i}/F_{i-1}\to 0,$
and computing the dimension, we obtain an inequality
$\dim(F_{i})^{\operatorname{GL}_{n}}\leq\dim(F_{i-1})^{\operatorname{GL}_{n}}+\dim(F_{i}/F_{i-1})^{\operatorname{GL}_{n}}$
and so
$\dim{\mathcal{R}}^{\operatorname{GL}_{n}}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{\operatorname{GL}_{n}}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$. Then the evaluation map
$f(x_{1},\ldots,x_{n})\mapsto[(v_{1},\ldots,v_{n})\in\operatorname{\mathbb{F}}_{q^{m}}^{n}\mapsto f(v_{1},\ldots,v_{n})\in\operatorname{\mathbb{F}}],$
induces an isomorphism of $\operatorname{\mathbb{F}}$-algebras
$\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}\xrightarrow{\cong}\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n},\operatorname{\mathbb{F}}\big{)}.$
This isomorphism is also $\operatorname{GL}_{n}$-equivariant where $\operatorname{GL}_{n}$ acts on the right-hand side via the embedding $\operatorname{GL}_{n}=\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q})\subset\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/\operatorname{GL}_{n},\operatorname{\mathbb{F}}\big{)}.$
The $\operatorname{\mathbb{F}}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}$ is thus equal to the cardinality of the orbit set $(\operatorname{\mathbb{F}}_{q^{m}})^{n}/\operatorname{GL}_{n}$. Finally this cardinality is equal to the sum $\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ where $\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ is the $q$-binomial coefficient which counts the number of $s$-dimensional subspaces of the $\operatorname{\mathbb{F}}_{q}$-vector space $\operatorname{\mathbb{F}}_{q^{m}}$.
In summary, we have shown that
$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $\operatorname{GL}_{n}$. The proposition is proved. ∎
We end this section with the following observation about our proposed basis $\mathcal{B}_{m}(\alpha)$.
Lemma 3.2.
For each composition $\alpha$ of $n$, the Hilbert-Poincaré series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
Proof.
Recall that
$C_{\alpha,m}(t)=\sum_{\beta\leq\alpha,|\beta|\leq m}t^{e(m,\alpha,\beta)}\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t},$
where
$e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})=(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})+q^{\beta_{1}}\sum_{i\geq 2}(\alpha_{i}-\beta_{i})(q^{m-\beta_{1}}-q^{\beta_{2}+\ldots+\beta_{i}}).$
Also, from [15, Section 7], we have the following identity for the $(q,t)$-multinomial coefficient:
$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\;\Biggl{(}\begin{bmatrix}m-\beta_{1}\\ \beta_{2},\ldots,\beta_{\ell},m-|\beta|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Thus if we write $\beta^{\prime}=(\beta_{2},\ldots,\beta_{\ell})$ and similarly for $\alpha^{\prime}$, then $C_{\alpha,m}(t)$ can be written as
$\sum_{\beta_{1}\leq\min(\alpha_{1},m)}t^{(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})}\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\biggl{(}\sum_{\beta^{\prime}\leq\alpha^{\prime},|\beta^{\prime}|\leq m-\beta_{1}}t^{e(m-\beta_{1},\beta^{\prime},\alpha^{\prime})}\begin{bmatrix}m-\beta_{1}\\ \beta^{\prime},m-\beta_{1}-|\beta^{\prime}|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Comparing this last formula with the recursive definition of $\mathcal{B}_{m}(\alpha)$, noting that $\delta_{s+1;m}^{\alpha_{1}-s}$ raises degree by $(\alpha_{1}-s)(q^{m}-q^{s})$, the Frobenius operator multiply the degree by $q$, and the Hilbert-Poincaré series for $\Delta^{m}_{s}$ is exactly $\begin{bmatrix}m\\ s\end{bmatrix}_{q,t}$, we see that the result follow easily by induction. ∎
The results in this section simplify our tasks considerably. Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}^{P_{\alpha}}(n)$, it suffices to verify that it is a generating set.
4. Proof of Proposition 1.6
As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. The proof of the Proposition 1.6 is done in two steps. Firstly, we use the relative transfer from the known invariant ring ${\mathcal{Q}}^{B}$ of the Borel subgroup $B$ to show that a bigger set $\mathcal{B}^{\prime}$, consisting of 3 families:
  1. (1)
    $\delta^{2}(1)=x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$,
  2. (2)
    $\delta_{2}(Q_{1,0}^{s})=y_{s}$, $0\leq s<[m]_{q}$,
  3. (3)
    ${\mathcal{D}}_{2}$,
is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. Then we show that to generate the invariant ring, it is possible to replace the Dickson algebra ${\mathcal{D}}_{2}$ in the third family by the smaller subspace $\Delta^{m}_{2}$.
Lemma 4.1.
The set $\mathcal{B}^{\prime}$ spans ${\mathcal{Q}}_{m}^{G}(2)$.
Proof.
In [6], we showed that an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families
  1. (i)
    $x_{1}^{q^{m}-1}x_{2}^{i(q-1)}$, $i\leq[m]_{q}$,
  2. (ii)
    $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$, $i_{1}<[m]_{q}$, $i_{2}\leq[m-1]_{q}$.
It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}^{B}(2)\to{\mathcal{Q}}_{m}^{G}(2)$ that it is onto, so that a generating set for ${\mathcal{Q}}^{G}(2)$ can be obtained from the image of $\operatorname{Tr}$.
By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}$ for any $0\leq s<[m]_{q}$:
$x_{1}^{q^{m}-1}x_{2}^{s(q-1)}=y_{0}x_{1}^{(s+1)(q-1)}-y_{s+1}.$
Since $y_{s}$ is already a $\operatorname{GL}_{2}$-invariant, this formula allows us to compute the transfer of the first family:
$\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{s(q-1)})=y_{0}\operatorname{Tr}(x_{1}^{(s+1)(q-1)})-y_{s+1}.$
Clearly, $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is again a Dickson polynomial in ${\mathcal{D}}_{2}$. By Proposition 2.5, the product of $y_{0}$ with a Dickson polynomial is either zero or another element in the $y$ family. Thus $\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})$ is either $-y_{s+1}$ or zero.
The transfer of the second family $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$ is clearly a Dickson polynomial. The lemma follows. ∎
The next step is to show that one only needs part of the Dickson algebra. The following observation about the Dickson monomials $Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}$ at the "edge" of $\Delta^{m}_{2}$ is crucial.
Proposition 4.2.
For each $10\leq i\leq m-$ 1, we have the following decomposition in ${\mathcal{Q}}_{m}(2)$:
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
Proof.
It is easy to see, by determinant manipulation, that $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})=y_{\frac{q^{i}-1}{q-1}}$ is a genuine Dickson polynomial. When $i=0$, the Dickson monomial $y_{0}$ must contains $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a montrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$. Furthermore, $y_{0}-Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is divisible by $Q_{2,0}$. Note that in degree $q^{m}-q$, $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is the only non-essential Dickson monomial. We have proved that
$y_{0}=Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}+\text{essential monomials divisible by $Q_{2,0}$}.$
For the general case, we make use of the identity $y_{m,\frac{q^{i}-1}{q-1}}=Q_{2,0}^{\frac{q^{i}-1}{q-1}}y_{m-i,0}^{q^{i}}$ where $y_{m,s}$ stands for $\delta_{2;m}(Q_{1,0}^{s})$. It follows that
$y_{\frac{q^{i}-1}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}+(fQ_{2,0})^{q^{i}}Q_{2,0}^{\frac{q^{i}-1}{q-1}},$
for some Dickson polynomial $f$. A Dickson monomial appearing in the second term of the sum above has the form $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ where $i_{2}\geq\frac{q^{i+1}-1}{q-1}$. Comparing the degree, we have $i_{1}\leq\frac{q^{m-1}-q^{i+1}}{q-1}$, hence $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ is essential. ∎
Remark 4.3.
It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another amusing formula:
$Q_{2,1}^{\frac{2(q^{m-1}-1)}{q-1}}=y_{\frac{q^{m}-q}{q-1}}=x_{1}^{q^{m}-q}x_{2}^{q^{m}-q}\quad\text{in ${\mathcal{Q}}_{m}(2)$}.$
We are now ready to finish the proof of Proposition 1.6.
Consider a Dickson monomial $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$ and suppose, by way of contradiction, that it is not in the span of $\mathcal{B}_{m}(2)$. Further, assume that it is the smallest monomial in grevlex order with this property. Since $Q_{2,0}^{k}=0$ if $k>[m-1]_{q}$, we must have $i_{2}\leq\frac{q^{m-1}-1}{q-1}=[m-1]_{q}$. Let $i\leq m-1$ be the unique integer such that $[i]_{q}\leq i_{2}<[i+1]_{q}$. Then $i_{1}\geq\frac{q^{m-1}-q^{i}}{q-1}$. It follows from Proposition 4.2 that
$Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}-Q_{2,1}^{i_{1}-\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{i_{2}-\frac{q^{i}-1}{q-1}}y_{\frac{q^{i}-1}{q-1}}$
is a sum of Dickson monomials in strictly smaller grevlex order than $(i_{1},i_{2})$, and at least one of them is again not in the span of $\mathcal{B}_{m}(2)$. We have a contradiction.
5. Proof of Proposition 1.7
It is straightforward to check that all elements in the six families listed in Proposition 1.7 are $P(2,1)$-invariants. It suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}^{P(2,1)}(3)$.
Suppose $f$ is a nonzero polynomials in ${\mathcal{Q}}_{m}^{P(2,1)}(3)$. If $f$ is considered as a polynomial in $x_{3}$, then its coefficients are $\operatorname{GL}_{2}$-invariant in the $2$ variables $x_{1},x_{2}$, which are classified in Section 4. If one of these coefficients has the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then the corresponding term belongs to the last 2 families in the list. By subtracting it if necessary, let us assume that this term does not occur in $f$.
Now suppose that the $x_{1}$ (and hence $x_{2}$ by symmetry) degree of $f$ is strictly less than $q^{m}-1$. Then $f$ is a linear combination of polynomials of the last two types in the list of Borel invariants in 2.5. In particular, the $x_{3}$-degree of $f$ must be divisible by $q^{2}-q$. Write $f$ as
$f=x_{3}^{aq(q-1)}f_{aq}+x_{3}^{(aq-1)(q-1)}f_{aq-1}+\ldots,$
where $f_{i}=f_{i}(x_{1},x_{2})\in{\mathcal{Q}}_{m}^{G}(2)$. If $f_{aq}$ contains $y_{t}$ for some $t\geq 0$, then we can subtract from $f$ an appropriate polynomial in the third or fourth family in the list 1.7. So we can assume that $f_{aq}$ belongs to $\Delta^{m}_{2}$. But this implies that this top degree monomial is a non-trivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. In particular, this shows that $a$ is divisible by $q$. Write $a=bq$. Then by subtracting from $f$ an appropriate linear combination of first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $G(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree. The result follows by induction on $x_{3}$-degree. ∎
Remark 5.1.
We can also prove that $\mathcal{B}_{m}(2,1)$ is linearly independent as follows. Consider these elements of $\mathcal{B}_{m}(2,1)$ as polynomials in $x_{3}$. The table below records the coefficient of the highest $x_{3}$-exponent.
$\mathcal{B}_{m}(2,1)$ $x_{3}$-degree Coefficient
(1) $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}Q_{3,2}^{i}$ $i(q^{3}-q^{2})<q^{m}-q^{2}$ $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(2) $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}\delta_{3}(1)$ $q^{m}-q^{2}$ $Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}}$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(3) $\delta_{2}(Q_{1,0}^{i_{1}}Q_{2,1}^{i})$ $i(q^{2}-q)<q^{m}-q$ $\delta_{2}(Q_{1,0}^{i_{1}})$, $i_{1}<[m]_{q}$
(4) $\delta_{2}(Q_{1,0}^{i_{1}}\delta_{2}(1))$ $q^{m}-q$ $\delta_{2}(Q_{1,0}^{i_{1}})$, $i_{1}<[m]_{q}$
(5) $\delta_{1}(\delta_{1}(Q_{1,0}^{i}))$ $i(q-1)<q^{m}-1$ $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$
(6) $\delta_{1}(\delta_{1}(\delta_{1}(1)))$ $q^{m}-1$ $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$
From the table, it is easy to see that the polynomials in $\mathcal{B}_{m}(2,1)$ are linearly independent.
6. Proof of Proposition 1.8
It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the smallest $x_{1}$-degree. Write
$f=x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})+x_{1}^{(i+1)(q-1)}g_{i+1}(x_{2},x_{3})+\ldots$
Note that all coefficient polynomials $g_{j}$ are $\operatorname{GL}_{2}$-invariant in $x_{2}$ and $x_{3}$. If $i(q-1)=q^{m}-1$, then $x_{1}^{q^{m}-1}g(x_{2},x_{3})$ is a linear combination of polynomials from the last 3 families of $\mathcal{B}_{m}(1,2)$. We can then subtract from $f$ this polynomial and can now assume that $i(q-1)\leq q^{m}-q$.
Since $f$ is a $B$-invariant, it is a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $Q_{1,0}^{j_{1}}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. By direct inspection, the lowest $x_{1}$-degree of these polynomials are $x_{1}^{j_{1}(q-1)}x_{2}^{q^{m}-q}x_{3}^{j_{2}q(q-1)}$ and $x_{1}^{j_{1}(q-1)}x_{2}^{j_{2}q(q-1)}Q_{2,1}^{qj_{3}}(x_{2},x_{3})$ respectively. In both cases, the exponent of $x_{2}$ and $x_{3}$ in the coefficient of the lowest $x_{1}$-degree is divisible by $q(q-1)$.
We conclude that $g_{i}(x_{2},x_{3})$ is a polynomial in which all exponents of $x_{2}$ and $x_{3}$ are divisible by $q(q-1)$. Since $g_{i}\in{\mathcal{Q}}_{m}^{G}(x_{2},x_{3})$, it must be a $q$-th power of a polynomial in ${\mathcal{Q}}_{m-1}(2)$ in the variables $x_{2},x_{3}$.
On the other hand, the lowest $x_{1}$-degree of the first 3 families of $\mathcal{B}_{m}(1,2)$ are
  1. (1)
    $x_{1}^{j_{1}(q-1)}[Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})]^{q}$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$;
  2. (2)
    $x_{1}^{j_{1}(q-1)}[\delta_{2;m-1}(Q_{1,0}^{j_{2}})]^{q}$, $j_{2}<[m-1]_{q}$;
  3. (3)
    $x_{1}^{j_{1}(q-1)}[x_{2}^{q^{m-1}-1}x_{3}^{q^{m-1}-1}]^{q}$.
Hence $x_{1}^{i(q-1)}g_{i}(x_{2},x_{3})$ is the lowest $x_{1}$-degree of a polynomial in the span of $\mathcal{B}_{m}(1,2)$. So by subtracting from $f$ an appropriate linear combination of polynomials in $\mathcal{B}_{m}(1,2)$, we obtain another $P(1,2)$-invariant polynomial whose smallest $x_{1}$-degree is strictly greater than that of $f$. By induction, the required result follows. ∎
Remark 6.1.
We can also prove directly $\mathcal{B}_{m}(1,2)$ is linearly independent. In fact, the generators can be distinguished by looking at the lowest $x_{1}$-degrees and the corresponding coefficients as given in the table below.
$\mathcal{B}_{m}(1,2)$ $x_{1}$-degree Coefficient
(1) $Q_{1,0}^{j_{1}}Q_{3,2}^{i_{1}}Q_{3,1}^{i_{2}}$ $j_{1}(q-1)<q^{m}-1$ $Q_{2,1}^{qi_{1}}(x_{2},x_{3})Q_{2,0}^{qi_{2}}(x_{2},x_{3})$, $(i_{1},i_{2})\in\Delta^{m-1}_{2}$
(2) $Q_{1,0}^{j_{1}}\delta_{3}(Q_{2,1}^{j_{2}})$ $j_{1}(q-1)<q^{m}-1$ $y_{m-1,j_{2}}^{q}(x_{2},x_{3})$, $j_{2}<[m-1]_{q}$
(3) $Q_{1,0}^{j_{1}}\delta_{2}(\delta_{2}(1))$ $j_{1}(q-1)<q^{m}-1$ $x_{2}^{q^{m}-q}x_{3}^{q^{m}-q}$
(4) $\delta_{1}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ $q^{m}-1$ $Q_{2,1}^{i_{1}}(x_{2},x_{3})Q_{2,0}^{i_{2}}(x_{2},x_{3})$, $(i_{1},i_{2})\in\Delta^{m}_{2}$
(5) $\delta_{1}(\delta_{2}(Q_{1,0}^{j_{1}}))$ $q^{m}-1$ $y_{j_{1}}(x_{2},x_{3}),j_{1}<[m]_{q}$
(6) $\delta_{1}(\delta_{1}(\delta_{1}(1)))$ $q^{m}-1$ $x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$
7. Proof of Proposition 1.9
We need to prove that $\mathcal{B}_{m}(3)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$. The proof’s strategy is the same as in the rank $2$ case, but the details are much more involved. The proof is divided into three steps. Using the transfer, we first construct a slightly bigger generating set $\mathcal{B}^{\prime}$ of the required form without restricting the Dickson polynomials:
$\mathcal{B}^{\prime}=\delta_{1}^{3}(\Delta^{m}_{0})\coprod\delta_{2}^{2}(\Delta^{m}_{1})\coprod\delta_{3}({\mathcal{D}}_{2})\coprod{\mathcal{D}}_{3}.$
Once this is done, we then show that $\delta_{3}({\mathcal{D}}_{2})$ is actually contained in the span of
$\delta_{1}^{3}(\Delta^{m}_{0})\coprod\delta_{2}^{2}(\Delta^{m}_{1})\coprod\delta_{3}(\Delta^{m}_{2}),$
and as a result, ${\mathcal{D}}_{2}$ can be replaced by the much smaller subset $\Delta^{m}_{2}$. Finally, we prove that when restricted to ${\mathcal{Q}}_{m}(3)$, the rank $3$ Dickson algebra belongs to the subspace spanned by $\mathcal{B}_{m}(3)$. This is proved by first showing that all the edge monomials of $\Delta^{m}_{3}$ are in this span. Thus the subspace spanned by $\mathcal{B}_{m}(3)$ is a ${\mathcal{D}}_{3}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. Since it contains $\Delta^{m}_{3}$, it will contain the entire ${\mathcal{D}}_{3}$.
Lemma 7.1.
The set $\mathcal{B}^{\prime}$ consisting of 4 families below is a generating set for ${\mathcal{Q}}_{m}^{G}(3)$.
  1. (1)
    $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{q^{m}-1}$.
  2. (2)
    $\delta_{2}^{2}(\Delta^{m}_{1})=\{a_{m,3,s}$, $0\leq s<[m]_{q}\}$.
  3. (3)
    $\delta_{3}({\mathcal{D}}_{2})$.
  4. (4)
    ${\mathcal{D}}_{3}$.
Proof.
We use the relative transfer from $P(1,2)$ to $G$. Among the six families in $\mathcal{B}_{m}(1,2)$, the last family is already a $G$-invariant. So are $\delta_{3}(Q_{2,1}^{j_{2}})$ and $\delta_{2}^{2}(1)$. On the other hand, $\operatorname{Tr}(x_{1}^{j_{1}(q-1)})$ is clearly an element of ${\mathcal{D}}_{3}$ and is a multiple of $Q_{3,0}$. It is known that
$Q_{3,0}\delta_{3}(Q_{2,1}^{j_{2}})=Q_{3,0}\delta_{2}^{2}(1)=0\quad\text{in ${\mathcal{Q}}_{m}$}.$
It follows that the transfer of the second and third family in $\mathcal{B}_{m}(1,2)$ contribute $G$-invariants of the form $\delta_{3}(Q_{2,1}^{j_{2}})$, $j_{2}<[m-1]_{q}$ and $\delta_{2}^{2}(1)$. The transfer of the 4th family clearly belongs to ${\mathcal{D}}_{3}$, so it remains to consider the 5th family.
We claim that if $j<[m]_{q}$, then
$\operatorname{Tr}_{P(1,2)}^{G}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=a_{m,3,j+1}.$
Indeed, consider $x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3})$ as a $B_{3}$-invariant and taking the transfer from $B_{3}$ to $P(2,1)$. In Section 4, we have already proved that
$\operatorname{Tr}_{B_{2}}^{\operatorname{GL}_{2}}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})=y_{a+1}(x_{1},x_{2}).$
Hence
$\operatorname{Tr}_{B_{3}}^{P(2,1)}(x_{1}^{q^{m}-1}y_{j}(x_{2},x_{3}))=\sum_{a=j}^{(q^{m}-q)/(q-1)}y_{a+1}(x_{1},x_{2})x_{3}^{q^{m}-q+j(a-1)-a(q-1)}.$
The last sum is exactly $a_{m,3,j+1}=\delta_{2}^{2}(Q_{1,0}^{j+1})$, which is already a $\operatorname{GL}_{3}$-invariant. By the transitivity property of the transfer, we have the required claim. The proof is finished. ∎
The next step is to show that in the third family, one need not use the entire ${\mathcal{D}}_{2}$. The following technical lemma will be used repeatedly.
Lemma 7.2.
Let $s,t$ be non-negative integers and $0\leq i<[m]_{q}$. Then the following equality holds in ${\mathcal{Q}}_{m}(3)$:
$\delta_{2}\big{(}V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2}(Q_{1,0}^{i})\big{)}=\begin{cases}0&\text{if $t\geq 1$ and $s>1$};\\ x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}&\text{if $t\geq 1$ and $s=1$};\\ (t+1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}&\text{if $t\geq 1$ and $s=0$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})&\text{if $t=0,s\leq 1$};\\ \delta_{2}^{2}(Q_{1,0}^{i+s})-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s+i-2)(q-1)}&\text{if $t=0,s\geq 2$}.\end{cases}$
Proof.
Recall that $\delta_{s+1}(f)$ is a polynomial if $f$ is a polynomial (possibly with higher than $s$ variables), which is $\operatorname{GL}_{s}$-invariant. Hence the expression under consideration is a polynomial. If $t\geq 1$, then by definition, $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{t(q-1)}\delta_{2;m}(Q_{1,0}^{i}))$ equals
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}-\\ x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}\bigg{]}\\ =x_{3}^{q^{m}}\;\frac{x_{2}^{q^{m}}x_{1}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{1}^{q^{m}}x_{2}^{(s+i)(q-1)+1}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}-\\ x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)+1}\;\frac{x_{1}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{1},x_{3})-x_{2}^{s(q-1)}V_{2}^{t(q-1)-1}(x_{2},x_{3})}{L_{2}(x_{1},x_{2})}.$
The first summand is clearly trivial in ${\mathcal{Q}}_{m}$ because the quotient involved is a polynomial. For the second summand, it is also trivial if $s>1$, and is $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-1)(q-1)}$ if $s=1$. If $s=0$, the result is
$-(t(q-1)-1)x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(qt+i-2)(q-1)}.$
Now suppose $t=0$. Then we have
$\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{2}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{i(q-1)}&x_{3}^{q^{m}}x_{1}^{i(q-1)}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\big{]}\\ =\frac{1}{L_{2}(x_{1},x_{2})}\big{[}x_{1}^{q^{m}}x_{2}\frac{\begin{vmatrix}x_{2}&x_{3}\\ x_{2}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{2}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}\frac{\begin{vmatrix}x_{1}&x_{3}\\ x_{1}^{q^{m}}x_{3}^{(s+i)(q-1)}&x_{3}^{q^{m}}x_{1}^{(s+i)(q-1)}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\big{]}+\\ \frac{x_{1}^{q^{m}}x_{2}^{q^{m}}x_{3}^{i(q-1)}}{L_{2}(x_{1},x_{2})}\Bigg{[}\frac{x_{2}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{2}^{q-1}-x_{3}^{q-1}}-\frac{x_{1}^{s(q-1)}-x_{3}^{s(q-1)}}{x_{1}^{q-1}-x_{3}^{q-1}}\Bigg{]}$
The first summand is $\delta_{2}^{2}(Q_{1,0}^{s+i})$. In ${\mathcal{Q}}_{m}(3)$, the second summand is trivial when $s\leq 1$ and equals $-x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{(s-2+i)(q-1)}$ if $s\geq 2$. ∎
We are now ready to show that
Lemma 7.3.
The subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by the following three families of elements
  1. (1)
    $\delta_{1}^{3}(\Delta^{m}_{0})$,
  2. (2)
    $\delta_{2}^{2}(\Delta^{m}_{1})$,
  3. (3)
    $\delta_{3}(\Delta^{m}_{2})$,
contains $\delta_{3}({\mathcal{D}}_{2})$.
Proof.
Suppose $f$ is a Dickson monomial in ${\mathcal{D}}_{2}$ but is not essential. Then it can be written in the form
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}Q_{2,1}^{a}Q_{2,0}^{b},$
for some $10\leq i\leq m-$ 1, $0\leq a,b$. Recall from 4.2 that
$Q_{2,1}^{\frac{q^{m-1}-q^{i}}{q-1}}Q_{2,0}^{\frac{q^{i}-1}{q-1}}=\delta_{2}(x^{q^{i}-1})+\text{essential monomials divisible by $Q_{2,0}^{\frac{q^{i+1}-1}{q-1}}$}.$
By repeating this process if necessary, we can write $f$ as a sum of essential monomials and Dickson polynomials of the form
$\delta_{2}(x^{q^{j}-1})Q_{2,1}^{a_{j}}Q_{2,0}^{b_{j}},\quad j\geq i.$
Thus it suffices to prove the lemma for $f=g\delta_{2}(x^{q^{i}-1})$, where $g\in{\mathcal{D}}_{2}$ and $10\leq i\leq m-$ 1.
To evaluate $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$, we first observe that by direct inspection, it can be simplified as a determinant:
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))=\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{2}^{q^{m}}&x_{3}^{q^{m}}\\ x_{1}^{q^{m}}g(x_{2},x_{3})&x_{2}^{q^{m}}g(x_{1},x_{3})&x_{3}^{q^{m}}g(x_{1},x_{2})\end{vmatrix}}{L_{3}(x_{1},x_{2},x_{3})}.$
Since
$L_{3}(x_{1},x_{2},x_{3})=$ $V_{3}(x_{1},x_{2},x_{3})L_{2}(x_{1},x_{2})$
$=$ $[V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}]V_{2}(x_{2},x_{3})L_{2}(x_{1},x_{2})$
$=$ $[V_{2}(x_{1},x_{3})^{q-1}-V_{2}(x_{1},x_{2})^{q-1}]V_{2}(x_{1},x_{3})L_{2}(x_{1},x_{2}),$
this determinant can be rewritten in the form:
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}h(x_{1},x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}h(x_{2},x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{L_{2}(x_{1},x_{3})}\bigg{]}.$
where
$h(x_{1},x_{2},x_{3})=\frac{g(x_{2},x_{3})-g(x_{2},x_{1})}{V_{2}(x_{2},x_{3})^{q-1}-V_{2}(x_{2},x_{1})^{q-1}}.$
This quotient is a symmetric polynomial in $V_{2}^{q-1}(x_{2},x_{3})$ and $V_{2}^{q-1}(x_{2},x_{1})$ with coefficient in $\mathbb{F}_{q}[x_{2}^{q-1}]$. Moreover, if one summand of this quotient is divisible by $V_{2}(x_{2},x_{1})^{q-1}$, then the common denominator $L_{2}(x_{1},x_{2})$ will be canceled out, and the result is zero in ${\mathcal{Q}}_{m}$.
Thus if we write $g$ in terms of the upper triangular invariants, $g=\sum_{s,t}V_{1}^{s(q-1)}V_{2}^{t(q-1)}$, then the determinant above equals in ${\mathcal{Q}}_{m}$ to a sum over all such pair $(s,t)$, $t\geq 1$ of
$\frac{1}{L_{2}(x_{1},x_{2})}\bigg{[}x_{1}^{q^{m}}x_{2}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{2},x_{3})\frac{\begin{vmatrix}x_{2}^{q^{i}}&x_{3}^{q^{i}}\\ x_{2}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{2},x_{3})}-x_{2}^{q^{m}}x_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}(x_{1},x_{3})\frac{\begin{vmatrix}x_{1}^{q^{i}}&x_{3}^{q^{i}}\\ x_{1}^{q^{m}}&x_{3}^{q^{m}}\end{vmatrix}}{V_{2}(x_{1},x_{3})}\bigg{]},$
which is $\delta_{2}(V_{1}^{s(q-1)}V_{2}^{(t-1)(q-1)}\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}}))$. We can now use the computation in Lemma 7.2 to see that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{2}(Q_{1,0}^{a}),a\geq 0)\}.$
On the other hand, $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})$ is a genuine Dickon polynomial, hence $\delta_{3}(g\delta_{2}(x^{q^{i}-1}))$ is a $\operatorname{GL}_{3}$-invariant in ${\mathcal{Q}}_{m}$. This shows that
$\delta_{3}(g\delta_{2}(x^{q^{i}-1}))\in\operatorname{Span}\{\delta_{2}^{2}(\Delta^{m}_{1}),\delta_{1}^{3}(1)\}.$
The third, and final step is to replace the full Dickson algebra ${\mathcal{D}}_{3}$ by $\Delta^{m}_{3}$. We first need some preliminary calculations. In the following, we write $\delta_{3;a}(i_{1},i_{2})$ for $\delta_{3;a}(Q_{2,1}^{i_{1}}Q_{2,0}^{i_{2}})$ to simplify the notation. Again $\delta_{3}$ will stand for $\delta_{3;m}$.
Lemma 7.4.
The following equalities hold:
  1. (1)
    $\delta_{3}(0,\frac{q^{i}-1}{q-1})=\delta_{3;m-i}^{q^{i}}(1)Q_{3,0}^{\frac{q^{i}-1}{q-1}}$ for all $i\geq 0$. In particular,
    $\delta_{3}(0,{\frac{q^{m-2}-1}{q-1}})=Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}\text{ and }\delta_{3}(0,\frac{q^{m-3}-1}{q-1})=Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}.$
  2. (2)
    $Q_{3,0}^{k}=0$ in ${\mathcal{Q}}_{m}$ for all $k>\frac{q^{m-2}-1}{q-1}$.
  3. (3)
    $\delta_{3}(q^{m-3},\frac{q^{m-3}-1}{q-1})=Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$.
  4. (4)
    In general, for $\lambda_{2}\geq\lambda_{3}\geq 0$, then
    $\delta_{3}\bigg{(}\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1}\bigg{)}=\delta_{3;m-\lambda_{3}}^{q^{\lambda_{3}}}\bigg{(}\frac{q^{\lambda_{2}-\lambda_{3}}-1}{q-1},0\bigg{)}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}.$
Proof.
The proof of these statements are quite similar, and essentially makes use of the Laplace expansion. We will prove the last statement only. Apply the Laplace expansion along the last row of the determinant in the numerator for $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})$, we obtain
$x_{1}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{2},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{2},x_{3})-x_{2}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{3})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3})+x_{3}^{q^{m}}Q_{2,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}(x_{1},x_{2})L_{2}^{q^{\lambda_{3}}}(x_{1},x_{3}).$
The result follows easily after writing this expression as a $q^{\lambda_{3}}$ power. ∎
Our second technical result plays a similar role as Proposition 4.2 in the case of 3 variables.
Proposition 7.5.
Suppose $20\leq\lambda\leq m-$ 2, then in ${\mathcal{Q}}_{m}$, $\delta_{3}(\frac{q^{\lambda}-1}{q-1},0)$ equals a Dickson polynomial in which the only non-essential summand which is not divisible by $Q_{3,0}$ is $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$.
Proof.
If $\lambda=0$, then $\delta_{3}(1)$ is a genuine Dickson polynomial in degree $q^{m}-q^{2}$. In this degree, there is only one Dickson monomial not in $\Delta^{m}_{3}$, namely $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$. If we set $x_{2}=x_{3}=0$, then both $Q_{3,1}$ and $Q_{3,0}$ becomes zero, $\delta_{3}(1)$ becomes $x_{1}^{q^{m}-q^{2}}$. Thus in the expansion of $\delta_{3}(1)$ into a sum of Dickson monomials, $Q_{3,2}^{\frac{q^{m-2}-1}{q-1}}$ must occur with non-zero coefficient. The lemma is proved for $\lambda=0$.
We now assume that $\lambda\geq 1$. Note that $\delta_{3}(Q_{2,1})=\frac{[0,2,m]}{[0,1,2]}$ is a genuine Dickson polynomial, hence in ${\mathcal{Q}}_{m}$,
$\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\delta_{3}(Q_{2,1})=Q_{3,2}^{\frac{q^{\lambda-1}-1}{q-1}}\frac{[0,2,m]}{[0,1,2]}.$
So $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ equals in ${\mathcal{Q}}_{m}$ to a Dickson polynomial and we are going to use this representation.
Setting $x_{3}=0$, $\delta_{3}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ is reduced to
$\frac{[\lambda+1,m]}{[1,2]}=y^{q^{\lambda+1}}_{m-\lambda-1,0}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}=Q_{2,1}^{\frac{q^{m-1}-q^{\lambda+1}}{q-1}}Q_{2,0}^{\frac{q^{\lambda+1}-q}{q-1}}+\text{others}.$
It follows that in this representation in ${\mathcal{Q}}_{m}$ of $\delta_{3;m}(Q_{2,1}^{\frac{q^{\lambda}-1}{q-1}})$ as a Dickson polynomial, $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,1}^{\frac{q^{\lambda-1}-1}{q-1}}$ must occur with nonzero coefficient. Moreover, any other monomials which involve $Q_{3,1}$ and $Q_{3,2}$ only must be divisible by $Q_{3,1}^{\frac{q^{\lambda}-1}{q-1}}$, and by degree reason, these monomials must be essential. ∎
The following corollary will be useful for induction.
Corollary 7.6.
Suppose $m-3\leq\lambda_{2}\leq\lambda_{3}\leq 0$ and $\lambda\leq m-2$.
  1. (1)
    $\delta_{3}(\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1},\frac{q^{\lambda_{3}}-1}{q-1})=Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$, where $\nu>\lambda_{3}$.
  2. (2)
    $\delta_{3}(\frac{q^{m-2}-q^{\lambda}}{q-1},\frac{q^{\lambda}-1}{q-1})=Q_{3,1}^{\frac{q^{m-2}-q^{\lambda}}{q-1}}Q_{3,0}^{\frac{q^{\lambda}-1}{q-1}}+\text{other terms},$
    where the other terms are either essential or non-essential and divisible by $Q_{3,0}^{\frac{q^{\nu}-1}{q-1}}$ where $\nu>\lambda$.
Proof.
This is just combining Lemma 7.4 and Proposition 7.5. ∎
Corollary 7.7.
The Dickson monomials at the edge of $\Delta^{m}_{3}$ listed below are also in the span of $\mathcal{B}_{m}(3)$:
  1. (1)
    $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$,
  2. (2)
    $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$, where $m-3\geq\lambda_{3}$,
  3. (3)
    $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$, where $m-3\geq\lambda_{2}\geq\lambda_{3}$.
Proof.
We have already seen from part (1) of Lemma 7.4 that $Q_{3,0}^{\frac{q^{m-2}-1}{q-1}}$ belongs to $\Delta^{m}_{3}$. For the other 2 families, we will prove by downward induction on $\lambda_{3}$.
If $\lambda_{3}=m-3$, then $\lambda_{2}=\lambda_{3}=m-3$. The edge monomials $Q_{3,1}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ and $Q_{3,2}^{q^{m-3}}Q_{3,0}^{\frac{q^{m-3}-1}{q-1}}$ belong to the span of $\mathcal{B}_{m}(3)$ by the last 2 parts of Lemma 7.4. We now suppose that the results have been verified for all edge elements corresponding to $\lambda_{3}+1$ and above. Consider the edge elements corresponding to $\lambda_{3}$:
  1. (1)
    $Q_{3,1}^{\frac{q^{m-2}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$,
  2. (2)
    $Q_{3,2}^{\frac{q^{m-2}-q^{\lambda_{2}}}{q-1}}Q_{3,1}^{\frac{q^{\lambda_{2}}-q^{\lambda_{3}}}{q-1}}Q_{3,0}^{\frac{q^{\lambda_{3}}-1}{q-1}}$.
Corollary 7.6 shows that these are in the image of $\delta_{3}(\Delta^{m}_{2})$ modulo those non-essential elements with strictly higher $\lambda_{3}$. By induction, these elements are already in the span of $\mathcal{B}_{m}(3)$. ∎
Now we are ready to finish the proof. Corollary 7.7 shows that the subspace of ${\mathcal{Q}}_{m}^{G}(3)$ spanned by $\mathcal{B}_{m}(3)$ is closed under the action of the Dickson algebra ${\mathcal{D}}_{3}$. Indeed, we have already computed the action of ${\mathcal{D}}_{3}$ on $\delta_{s+1}^{3-s}(\Delta^{m}_{s})$ for $0\leq s\leq 2$ in 2.5, and the product of $Q_{3,i}$ with a monomial in $\Delta^{m}_{3}$ is clearly an edge monomial, which is in the span of $\mathcal{B}_{m}(3)$ by the above demonstration. The proof is finished.
8. A filtration of submodules
For each positive integer $n$, $0\leq k\leq n$, let $\mathcal{F}_{n,k}$ denote the $\mathbb{F}_{q}$-subspace of ${\mathcal{Q}}^{G}_{m}(n)$:
$\mathcal{F}_{n,k}=\operatorname{Span}(\delta_{s+1}^{n-s}(f)\colon f\in\Delta^{m}_{s},0\leq s\leq\min(m,k)).$
For $n=3$, we have shown that this increasing filtration of subspaces is exhaustive. The calculation in previous sections shows that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{G}(3)$ is natural in the sense that the resulting filtration is a filtration of ${\mathcal{D}}_{3}$-submodules. Moreover, the action of the Dickson invariants is very similar to the pattern discussed in [10] about the structure of the cofixed space $S_{G}$ as a $S^{G}$-module. Indeed, we can easily check that $Q_{3,0}$ annihilates $\mathcal{F}_{3,2}$; $Q_{3,0}$ and $Q_{3,1}$ annihilate $\mathcal{F}_{3,1}$; and $Q_{3,0},Q_{3,1},Q_{3,2}$ annihilate $\mathcal{F}_{3,0}$. A similar pattern also occurs in the case $n=2$.
The aim of this final section is to demonstrate Theorem 1.11. We will show the filtration $\mathcal{F}_{n,*}$ also behaves well with respect to the action of the mod $q$ Steenrod algebra $\mathcal{A}$, at least in the case $n\leq 3$.
For background on the mod $q$ Steenrod algebra and its role in modular invariant theory, we refer to Larry Smith’s article [16], which treats the Steenrod algebra completely algebraically over a general finite field.
Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is a $\mathbb{F}_{q}$-linear map satisfying the following conditions:
  1. (1)
    The unstable condition: $\mathcal{P}^{i}(f)=f^{q}$ if $i=\deg(f)$ and $\mathcal{P}^{i}(f)=0$ if $i>\deg(f)$.
  2. (2)
    The Cartan formulae: $\mathcal{P}^{i}(fg)=\sum_{a+b=i}\mathcal{P}^{a}(f)\mathcal{P}^{b}(g)$.
Moreover, $\mathcal{P}^{0}$ acts as the identity. Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which $x_{i}$ are given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae. More concretely, we have
$\mathcal{P}^{i}(v^{j})=\binom{j}{i}v^{j+i(q-1)},$
for any linear polynomial $v$. Here the binomial coefficient is taken modulo $p$. Thus, if $j$ is a $q$-power, then $\mathcal{P}^{i}(v^{j})\neq 0$ iff $i=0$ (in which case $\mathcal{P}^{0}$ is the identity operator) or $i=j$ (in which case $\mathcal{P}^{i}(v^{i})=v^{qi}$ is the Frobenius operator).
The Steenrod algebra action and the $G$ action on $S$ commutes. Furthermore, it is clear that $I_{m}$ is stable under the action of the Steenrod operations, so is ${\mathcal{Q}}_{m}=S/I_{m}$. It follows that a Steenrod reduced power yields a map from ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ to itself. In other words, ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is an $\mathcal{A}$-module.
The proof of Theorem 1.11 uses several lemmas. We begin with a general formula describing how the Steenrod operation interacts with the delta operator.
Lemma 8.1.
If $f$ is a polynomial which is $\operatorname{GL}_{1}$-invariant in the first variable and $1\leq k\leq\deg(\delta_{2}(f))=q^{m}-q+\deg(f)$, then
$\mathcal{P}^{k}(\delta_{2}f)+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}f)+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}f)=\\ \delta_{2}(Q_{1,0}\mathcal{P}^{k-1}f)+\delta_{2}(\mathcal{P}^{k}f)+\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1-q^{m}}f)+\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f).$
Moreover, the last two summands on the right-hand side vanish in ${\mathcal{Q}}_{m}$.
Proof.
The condition imposed on $f$ assures that $\delta_{2}(f)$ as well as the other terms in this lemma are polynomial. The equation above is then obtained by applying the Cartan formula for the product $L_{2}\delta_{2}(f)$ and then dividing both sides by $L_{2}$. Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all $k\geq 1$. ∎
Corollary 8.2.
For each $0\leq k\leq 2$, $\mathcal{F}_{2,k}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(2)$ over the Steenrod algebra.
Proof.
We need to compute the action of the Steenrod operation on a typical element $\delta_{2}(f)$ where $f=Q_{1,0}^{a}$, $a<[m]_{q}$. In the rank $2$ case, the last two summands on the right-hand side of the formula given in Lemma 8.1 vanish in ${\mathcal{Q}}_{m}$. Indeed, one check by direct inspection that the degree of $\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1}f)$ is strictly greater than $2(q^{m}-1)$ for all $q$ and $m$. For $\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f)$, its degree is not greater than $2(q^{m}-1)$ only when $q=2$, $k\geq q^{m}$ and $\deg f=0$, but in this case $k>q^{m}-q+\deg(f)$.
Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.5. ∎
Corollary 8.3.
$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(3)$ over the Steenrod algebra.
Proof.
We need to show that for each $k\geq 1$ and $a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$.
Apply the formula provided in Lemma 8.1 for $f=\delta_{2}Q_{1,0}^{a}$. Again, the last two summands are trivial in ${\mathcal{Q}}_{m}$ because of the instability condition and degree calculation. We then obtain a recursive equation in ${\mathcal{Q}}_{m}$:
$\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}Q_{1,0}^{a})+\delta_{2}(\mathcal{P}^{k}\delta_{2}Q_{1,0}^{a}).$
We now proceed by double induction on $k$ that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ belongs to $\mathcal{F}_{3,1}$. We already know from Corolarry 8.2 that $\mathcal{P}^{k-1}(\delta_{2}Q_{1,0}^{a})$ and $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{a})$ belongs to the span of $\delta_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}(1)$. By Lemma 7.2, the two summands on the right-hand side in the equation above belong to the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$.
Also by induction, we see that $Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=0$ in ${\mathcal{Q}}_{m}$ and $Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$. We conclude that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$, but since it is a $\operatorname{GL}_{3}$-invariant, it is spanned by $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{3}(1)$. ∎
We now consider the interaction between the Steenrod powers and the operator $\delta_{3}$.
Lemma 8.4.
If $f\in{\mathcal{D}}_{2}$ and $k\geq 1$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)=\frac{\mathcal{P}^{k}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ x_{1}^{q^{m}}f(x_{2},x_{3})&x_{2}^{q^{m}}f(x_{1},x_{3})&x_{3}^{q^{m}}f(x_{1},x_{2})\end{vmatrix}}{L_{3}}.$
Moreover, the right-hand side equals $\delta_{3;m}(F)$ or $\delta_{3;m+1}(F)$ where $F$ is a certain linear combination of polynomials of the form $g\mathcal{P}^{i}f$ where $g\in{\mathcal{D}}_{2}$ and $0\leq i\leq k$.
Proof.
The above equality is a simple application of the Cartan formula for $\mathcal{P}^{k}(L_{2}\delta_{2;m}(f))$ (or $\mathcal{P}^{k}(L_{3}\delta_{3}f)$). For the right-hand side, note that because of the instability condition, we have $\mathcal{P}^{i}(x_{1})$ is either $x_{1}$ or $x_{1}^{q}$, depending on whether $i=0$ or $i=1$, and there is a similar formula for other entries in the determinant.
In the first case, $\delta_{2;m+1}(g\mathcal{P}^{i}f)$ is trivial in ${\mathcal{Q}}_{m}$ unless $q=2$ and $g\mathcal{P}^{i}f=1$. In this case, $\delta_{2;m+1}(1)=x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}$ because of degree reason. We also have $g=f=1$, and $k=2^{m}$, but then $\delta_{2}f$ has degree $2^{m}-2<2^{m}$.
Finally, observe that any determinant of the form
$\begin{vmatrix}x_{1}^{q^{a}}&x_{2}^{q^{a}}&x_{3}^{q^{a}}\\ x_{1}^{q^{b}}&x_{2}^{q^{b}}&x_{3}^{q^{b}}\\ x_{1}^{q^{c}}\mathcal{P}^{d}f(x_{2},x_{3})&x_{2}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{3})&x_{3}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{2})\end{vmatrix}$
can be written as
$L_{3}\delta_{3;c}\bigg{(}\frac{[a,b]}{[0,1]}\mathcal{P}^{d}f\bigg{)}.$
where $\frac{[a,b]}{[0,1]}\in\mathcal{D}_{2}$. Note that $c$ is either $m$ or $m+1$. ∎
It is also clear that $\mathcal{P}^{i}L_{3}$ is again a matrix whose entries are $q$-powers, so it is always a multiple of $L_{3}$ and the quotient $\frac{\mathcal{P}^{i}L_{3}}{L_{3}}$ is a Dickson polynomial in ${\mathcal{D}}_{3}$. The next lemma discusses how to deal with the case when $c=m+1$.
Lemma 8.5.
If $h\in{\mathcal{D}}_{2}$, then the following holds in ${\mathcal{Q}}_{m}$,
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $q>3$ or $q=3$ and $\deg(h)>0$;}\\ \delta_{3;m}(Q_{2,1}^{q^{m-1}})&\text{if $q=3$ and $h=1$;}\\ \in\operatorname{Span}\{\delta_{2;m}^{2}(\Delta^{m}_{1}),\delta_{1;m}^{3}(1)\}&\text{if $q=2$.}\\ \end{cases}$
Proof.
The degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$. If $q>3$, then this is strictly greater than $3(q^{m}-1)$, and we immediately have $\delta_{3;m+1}(h)\equiv 0$ in this case.
From the fundamental equation, we know that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,$
and so
$\delta_{3;m+1}(h)-Q_{3,2}^{q^{m-2}}\delta_{3;m}(h)+Q_{3,1}^{q^{m-2}}\delta_{3;m-1}(h)-Q_{3,0}^{q^{m-2}}\delta_{3;m-2}(h)=0.$
If $q=3$, then $q^{m+1}-q^{2}+\deg(h)\leq 3(q^{m}-1)$ implies $\deg(h)\leq q^{2}-3=6$. Since $h\in{\mathcal{D}}_{2}$, it follows that up to a scalar, $h=1$ or $h=Q_{2,1}$.
In the case where $h=Q_{2,1}$, $\delta_{3;m+1}(Q_{2,1})$ has degree $33^{m+1}-$ 3 so that it is either zero or the top element $x_{1}^{3^{m}-1}x_{2}^{3^{m}-1}x_{3}^{3^{m}-1}$. We claim that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. Indeed, when $q=3$, then by direct computation, we have $Q_{3,0}=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}$. It follows that in ${\mathcal{Q}}_{m}$, we have
$\delta_{3;m+1}(Q_{2,1})=Q_{3,2}^{q^{m-2}}\delta_{3;m}(Q_{2,1}).$
On the other hand, $Q_{3,2}=x_{1}^{6}x_{2}^{6}x_{3}^{6}$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,2}^{q^{m-2}}=x_{1}^{2.3^{m-1}}x_{2}^{2.3^{m-1}}x_{3}^{2.3^{m-1}}$ in ${\mathcal{Q}}_{m}$. So if we can show that $\delta_{3;m}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m-1}$, then it follows immediately that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. The claim can now be easily verified by checking that $\delta_{3;3}(Q_{2,1})=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$.
Suppose now $h=1$. In this case, we have
$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}),$
as required. Here the second equality follows from 2.5.
Now consider the case $q=2$. We still have $Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$, but now
$Q_{3,1}^{2^{m-2}}=x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}.$
By direct inspection, we have $\delta_{3;m+1}(h)=Q_{3,2}^{2^{m-2}}\delta_{3;m}(h)-Q_{3,1}^{2^{m-2}}\delta_{3;m-1}(h)$ equals in ${\mathcal{Q}}_{m}$ to
$\frac{1}{L_{3}}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{2}&x_{2}^{2}&x_{3}^{2}\\ x_{1}^{2^{m}}h(x_{2},x_{3})(x_{2}^{2^{m}}+x_{3}^{2^{m}})&x_{2}^{2^{m}}h(x_{1},x_{3})(x_{1}^{2^{m}}+x_{3}^{2^{m}})&x_{3}^{2^{m}}h(x_{1},x_{2})(x_{1}^{2^{m}}+x_{2}^{2^{m}})\end{vmatrix}.$
It is straightforward to verify that
$x_{1}^{2^{m}}+x_{2}^{2^{m}}=Q_{2,1}\delta_{2;m}(1)-\delta_{2;m}(Q_{1,0}^{2}),$
so the expression above becomes
$\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
Now we can use the argument as in the proof of Lemma 7.3 to conclude. ∎
Remark 8.6.
With a little more work, using 7.2, it can be showed that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have an explicit formula:
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$,}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$}.\end{cases}$
We leave the details for the interested reader.
Now we can finish the proof of Theorem 1.11. It remains to show that $\mathcal{F}_{3,1}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. From Lemma 8.5 we see that if $f\in{\mathcal{D}}_{2}$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$
belongs to $\mathcal{F}_{3,1}$. By induction, so is $\mathcal{P}^{i}(\delta_{3}f)$ for each $i<k$ and from Proposition 2.5, $\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,1}$.